OpenStudy (anonymous):

Find all real number solutions for: x^3-5x^2+8x-4

OpenStudy (schrodingers_cat):

First use the rational root test. Namely, find all possible combinations of q/p where q is the factors of the constant and p is the factors of the leading coefficient. You will end up with ±1/1, ±2/1, ±2/1, ±4/1 Plugging each of these values into the polynomial you will find that 1/1 is rational zero of the function. Then you can use the factor theorem and divide the polynomial by qx-p and solve for the rest of the zeros. So you will get, \[x^3-5x^2+8x-4/(x-1) = x^2 -4x +4\] You can get this result by synthetic division or long division Factoring this result you get x^2 -4x +4 = (x-2)^2 So, in all you get \[x^3-5x^2+8x-4 =(x-1)(x-2)^2\] Hope this helps :)

OpenStudy (anonymous):

thank you